# The quantum-refractive evolution of polarization states in pulsar emission

### v1

@@ 2.1 Evolution equations of Stokes vector
$\hat{\Omega}_{2}$ $\approx 2\eta B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (19)
where $\eta\equiv\alpha_{\mathrm{e}}/\left(30\pi B_{\mathrm{c}}^{2}\right)$ and
@@
with $B_{\mathrm{max}}$ being the maximum magnetic field strength at the polar cap [^6] From equation (5) $B_{\mathrm{max}}=\left|B\left(r=r_{*},\theta=\alpha\right)\right|$. and $r_{*}$ being the neutron star radius ($\approx 10^{6}\,\mathrm{cm}$), and $\hat{n}_{y\left[0\right]}$, $\cos\vartheta$ and $\sin\vartheta$ are given by equations (11), (16) and (17), respectively.
To facilitate solving the evolution equations (1) in the next Subsection, we substitute the following identities,
$\cos\left(\theta-\alpha\right)=\frac{\sin\alpha x+\cos\alpha z}{\left(x^{2}+z^{2}\right)^{1/2}},~{}~{}\sin\left(\theta-\alpha\right)=\frac{\cos\alpha x-\sin\alpha z}{\left(x^{2}+z^{2}\right)^{1/2}},~{}~{}\sin\left(\theta-\theta_{\mathrm{o}}\right)=\frac{\cos\theta_{\mathrm{o}}x-\sin\theta_{\mathrm{o}}z}{\left(x^{2}+z^{2}\right)^{1/2}},$ (21)
@@
$x=\hat{n}_{x\left[0\right]}s+r_{\mathrm{o}}\sin\theta_{\mathrm{o}},~{}~{}z=\hat{n}_{z\left[0\right]}s+r_{\mathrm{o}}\cos\theta_{\mathrm{o}}$ (22)
into (16) and (17). Then our solutions for the Stokes vector $\mathbf{S}$ will be parametrised by $s$.
2.2 Solving evolution equations
From equation (1) we write down a system of first-order ordinary differential equations to solve:
@@ 2.2 Solving evolution equations
However, in case $\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\ll 1$, one can obtain a solution to equation (1) via perturbation:
@@
where $\delta\mathbf{S}_{\left[1\right]}$ means the leading order quantum correction to the unperturbed (initial) Stokes vector $\mathbf{S}_{\left[0\right]}$. Here the correction can be treated as the leading order perturbation with $\alpha_{\mathrm{e}}/\left(30\pi\right)\left(B/B_{\mathrm{c}}\right)^{2}\sim 10^{-5}\left(B/B_{\mathrm{c}}\right)^{2}$ being a perturbation parameter. Upon inspection of equations (18) and (19) for equation (26), we can further write down our solution in terms of its components:
$S_{1}$ $\approx S_{1\left[0\right]}+2k\eta S_{3\left[0\right]}\int B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]\mathrm{d}s,$ (27)
@@
$S_{3}$ $\approx S_{3\left[0\right]}-k\eta\left\{S_{2\left[0\right]}\int B^{2}\sin^{4}\vartheta\mathrm{d}s+2S_{1\left[0\right]}\int B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]\mathrm{d}s\right\}.$ (29)
@@ 2.2.1 Examples
We consider X-ray emissions from three neutron stars, (i) one with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), (ii) another with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), (iii) the third with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), and assume for all the emissions, $r_{\mathrm{o}}=2r_{*}\approx 2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\eta\approx 3.97\times 10^{-32}$ and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ ($k\approx 2.0958\times 10^{8}\,\mathrm{cm^{-1}}$). These stars belong to ‘rotation-powered pulsars’ (RPPs) (Pavlov et al. (2013)). [^7] RPPs refer to neutron stars whose radiation is powered by loss of their rotation energy, via creation and acceleration of $e^{+}e^{-}$ pairs in the strong magnetic field, $B_{\mathrm{max}}\sim 10^{11}-10^{13}\,\mathrm{G}$. The number of detected RPPs are known to be about $\sim 2000$ in radio, $\sim 10$ in optical (including NIR and UV), $\sim 100$ in X-ray and $\sim 100$ in gamma-ray emissions (Pavlov et al. (2013)). In Fig. 2 the three RPPs chosen from the X-ray group are encircled, (i) the one in orange, (ii) another in cyan, (iii) the third in green. Given the X-ray emissions from these, we solve the evolution equations (1) for the following cases.
Figure 2: The population of detected rotation-powered pulsars (RPPs) plotted against their rotation period. RPPs detected in X-rays and gamma-rays are represented by red dots and blue stars, respectively. Three RPPs chosen from the X-ray group, (i) one with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), (ii) another with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), (iii) the third with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), are encircled in orange, cyan and green colours, respectively. (Credit: Pavlov et al. (2013), reproduced with modifications.)
@@
With $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), we obtain numerical solutions through equations (23)-(25) (in solid lines) or perturbative solutions by means of equations (27)-(29) (in dashed lines), as shown in Figs. 3a and 3b, given the initial Stokes vectors $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)=\left(1,0,0\right)$ and $\left(0.8,0,0.6\right)$, respectively. The perturbative solutions agree well with numerical ones as $\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\sim\left|k\hat{\Omega}_{1,2\left(\mathrm{min}\right)}s_{\dagger 1,2}\right|\sim 10^{-2}\ll 1$, where $\hat{\Omega}_{1,2\left(\mathrm{min}\right)}\equiv\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)$, evaluated from the birefringent functions given by (18), (19), with $s_{\dagger 1,2}$ being the extremum, i.e., $\mathrm{d}\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)/\mathrm{d}s=0$. On the Poincaré sphere, our solutions are represented by the magenta and light blue loci in Fig. 6a, corresponding to Figs. 3a and 3b, respectively. The loci imply a fraction of an oscillation for the polarization evolution, as is confirmed later by the approximate analytical solutions in Section 2.2.2.
@@
Figure 3: Corresponding to Example (i): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$).
@@
With $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), we obtain numerical solutions through equations (23)-(25), as shown in Figs. 4a and 4b, given the initial Stokes vectors $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)=\left(1,0,0\right)$ and $\left(0.8,0,0.6\right)$, respectively. On the Poincaré sphere, our solutions are represented by the magenta and light blue loci in Fig. 6b, corresponding to Figs. 4a and 4b, respectively. The loci imply about half an oscillation for the polarization evolution, as is confirmed later by the approximate analytical solutions in Section 2.2.2.
@@
Figure 4: Corresponding to Example (ii): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$).
@@
With $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), we obtain numerical solutions through equations (23)-(25), as shown in Figs. 5a and 5b, given the initial Stokes vectors $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)=\left(1,0,0\right)$ and $\left(0.8,0,0.6\right)$, respectively. On the Poincaré sphere, our solutions are represented by the magenta and light blue loci in Fig. 6c, corresponding to Figs. 5a and 5b, respectively. The loci imply multiple oscillations for the polarization evolution, as is confirmed later by the approximate analytical solutions in Section 2.2.2.
@@
Figure 5: Corresponding to Example (iii): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$).
(a) Corresponding to Example (i):
a fraction of an oscillation
(b) Corresponding to Example (ii):
about half an oscillation
(c) Corresponding to Example (iii):
multiple oscillations
Figure 6: Representations of the Stokes vectors on the Poincaré sphere. The loci imply patterns of the polarization evolution in terms of oscillation.

### v3

@@ 2.1 Evolution equations of Stokes vector
$\hat{\Omega}_{2}$ $\approx-2\eta B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha\,s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]$
$\hskip 12.0pt+\mathcal{O}\left(\phi^{2},\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)^{2},\phi\left(\Omega_{\mathrm{p}}r_{\mathrm{o}}/c\right)\right),$ (19)
where $\eta\equiv\alpha_{\mathrm{e}}/\left(30\pi B_{\mathrm{c}}^{2}\right)$666 $\eta=\eta_{2}-\eta_{1}$, where $\eta_{1}$ and $\eta_{2}$ are parameters defined via $\eta_{1}/4=\eta_{2}/7=\alpha_{\mathrm{e}}/\left(90\pi B_{\mathrm{c}}^{2}\right)\sim 10^{-31}\,\mathrm{g}^{-1}\,\mathrm{cm}\,\mathrm{s}^{2}$, from the post-Maxwellian Lagrangian $\mathcal{L}_{\mathrm{PM}}=-\left(\mathbf{B}^{2}-\mathbf{E}^{2}\right)/2+\eta_{1}\left(\mathbf{B}^{2}-\mathbf{E}^{2}\right)^{2}/4+\eta_{2}\left(\mathbf{E}\cdot\mathbf{B}\right)^{2}$ Euler1935Ueber . and
@@
with $B_{\mathrm{max}}$ being the maximum magnetic field intensity at the polar cap [^7] From Eq. (5) $B_{\mathrm{max}}=\left|B\left(r=r_{*},\theta=\alpha\right)\right|$. and $r_{*}$ being the neutron star radius ($\approx 10^{6}\,\mathrm{cm}$), and $\hat{n}_{y\left[0\right]}$, $\cos\vartheta$ and $\sin\vartheta$ are given by Eqs. (11), (16) and (17), respectively.
To facilitate solving the evolution equation (1) in the next subsection, we substitute the following identities,
$\cos\left(\theta-\alpha\right)=\frac{\sin\alpha\,x+\cos\alpha\,z}{\left(x^{2}+z^{2}\right)^{1/2}},~{}\sin\left(\theta-\alpha\right)=\frac{\cos\alpha\,x-\sin\alpha\,z}{\left(x^{2}+z^{2}\right)^{1/2}},$
$\sin\left(\theta-\theta_{\mathrm{o}}\right)=\frac{\cos\theta_{\mathrm{o}}\,x-\sin\theta_{\mathrm{o}}\,z}{\left(x^{2}+z^{2}\right)^{1/2}},$ (21)
@@
$x=\hat{n}_{x\left[0\right]}s+r_{\mathrm{o}}\sin\theta_{\mathrm{o}},~{}z=\hat{n}_{z\left[0\right]}s+r_{\mathrm{o}}\cos\theta_{\mathrm{o}}$ (22)
into (16) and (17). Then our solutions for the Stokes vector $\mathbf{S}$ will be parameterized solely by $s$.
2.2 Solving the evolution equations
From Eq. (1) we write down a system of first-order ordinary differential equations to solve:
@@ 2.2 Solving the evolution equations
However, in case $\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\ll 1$, one can obtain a solution to Eq. (1) via perturbation:
@@
where $\delta\mathbf{S}_{\left[1\right]}$ means the leading order quantum correction to the unperturbed (initial) Stokes vector $\mathbf{S}_{\left[0\right]}$. Here the correction can be treated as the leading order perturbation with $\alpha_{\mathrm{e}}/\left(30\pi\right)\left(B/B_{\mathrm{c}}\right)^{2}\sim 10^{-5}\left(B/B_{\mathrm{c}}\right)^{2}$ being a perturbation parameter. Upon inspection of Eqs. (18) and (19) for Eq. (26), we can further write down our solution in terms of its components:
$S_{1}$ $\approx S_{1\left[0\right]}-2k\eta S_{3\left[0\right]}\int B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha\,s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]\mathrm{d}s,$ (27)
@@
$S_{3}$ $\approx S_{3\left[0\right]}+k\eta\left\{2S_{1\left[0\right]}\int B^{2}\sin^{3}\vartheta\left[\left(\cos\vartheta+\sin\vartheta\right)\hat{n}_{y\left[0\right]}-\frac{\Omega_{\mathrm{p}}\sin\alpha\,s}{c\left(3\cos^{2}\left(\theta-\alpha\right)+1\right)^{1/2}}\right]\mathrm{d}s\right.$
$\!\left.\left.\hskip 12.0pt-S_{2\left[0\right]}\int B^{2}\sin^{4}\vartheta\mathrm{d}s\right._{\,{}_{\,{}_{\,}}}^{\,{}^{\,{}^{\,}}}\!\!\!\!\!\right\}.$ (29)
@@ 2.2.1 Examples
We consider X-ray emissions from three neutron stars: (i) one with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), (ii) another with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), (iii) the third with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$). For all three, we assume $r_{\mathrm{o}}=2r_{*}\approx 2\times 10^{6}\,\mathrm{cm}$, $\theta_{\mathrm{o}}=60^{\circ}$, $\alpha=45^{\circ}$, $\eta\approx 3.97\times 10^{-32}$ and $\omega\approx 2\pi\times 10^{18}\,\mathrm{Hz}$ ($k\approx 2.0958\times 10^{8}\,\mathrm{cm^{-1}}$). [^8] For the emission location $\left(r_{\mathrm{o}},\theta_{\mathrm{o}}\right)$ and the inclination angle $\alpha$ are given the same values for the three stars; the values are not based on actual observations. This is intended for comparing the QED effects from the three different sources under the same conditions. These stars belong to ‘rotation-powered pulsars’ (RPPs) Pavlov2013 . [^9] RPPs refer to neutron stars whose radiation is powered by loss of their rotation energy, via creation and acceleration of $e^{+}e^{-}$ pairs in the strong magnetic field, $B_{\mathrm{max}}\sim 10^{11}-10^{13}\,\mathrm{G}$. The number of detected RPPs are known to be about $\sim 4000$ in radio, $\sim 10$ in optical (including NIR and UV), $\sim 100$ in X-ray and $\sim 300$ in gamma-ray emissions Pavlov2013 ; ANTF2024 ; Smith_2023 . In Fig. 2 the three RPPs chosen from the X-ray group are encircled: (i) the one in orange, (ii) another in cyan, (iii) the third in green. Given the X-ray emissions from these, we solve the evolution equations (1) for the following cases.
Figure 2: The population of detected rotation-powered pulsars (RPPs) plotted against their rotation period. RPPs detected in X-rays and gamma-rays are represented by red dots and blue stars, respectively. Three RPPs chosen from the X-ray group, (i) one with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), (ii) another with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), (iii) the third with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), are encircled in orange, cyan and green colors, respectively. (Credit: Pavlov2013 , reproduced with modifications.)
@@
With $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$), we obtain numerical solutions of Eqs. (23)-(25) (in solid lines) or perturbative solutions by means of Eqs. (27)-(29) (in dashed lines), as shown in Figs. 3a and 3b, given the initial Stokes vectors $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)=\left(1,0,0\right)$ and $\left(0.8,0,0.6\right)$, respectively. The perturbative solutions agree well with numerical ones as $\left|k\hat{\Omega}_{1,2}\left(s\right)s\right|_{\mathrm{max}}\sim\left|k\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)s_{\dagger 1,2}\right|\sim 10^{-2}\ll 1$, where $s_{\dagger 1,2}$ is the extremum, i.e., $\mathrm{d}\hat{\Omega}_{1,2}\left(s_{\dagger 1,2}\right)/\mathrm{d}s=0$. On the Poincaré sphere, our solutions are represented by the magenta and light blue loci in Fig. 6a, corresponding to Figs. 3a and 3b, respectively. The loci imply a fraction of an oscillation for the polarization evolution, as is confirmed later by the approximate analytical solutions in Sect. 2.2.2.
@@
Figure 3: For Example (i): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 392.7\,\mathrm{Hz}$ ($P\approx 0.016\,\mathrm{s}$); the subscript $\left(\mathrm{p}\right)$ stands for ‘perturbative’.
@@
With $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$), we obtain numerical solutions of Eqs. (23)-(25), as shown in Figs. 4a and 4b, given the initial Stokes vectors $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)=\left(1,0,0\right)$ and $\left(0.8,0,0.6\right)$, respectively. On the Poincaré sphere, our solutions are represented by the magenta and light blue loci in Fig. 6b, corresponding to Figs. 4a and 4b, respectively. The loci imply about half an oscillation for the polarization evolution, as is confirmed later by the approximate analytical solutions in Sect. 2.2.2.
@@
Figure 4: For Example (ii): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 5.6\times 10^{12}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 22.28\,\mathrm{Hz}$ ($P\approx 0.282\,\mathrm{s}$).
@@
With $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$), we obtain numerical solutions of Eqs. (23)-(25), as shown in Figs. 5a and 5b, given the initial Stokes vectors $\mathbf{S}\left(0\right)=\left(S_{1}\left(0\right),S_{2}\left(0\right),S_{3}\left(0\right)\right)=\left(1,0,0\right)$ and $\left(0.8,0,0.6\right)$, respectively. On the Poincaré sphere, our solutions are represented by the magenta and light blue loci in Fig. 6c, corresponding to Figs. 5a and 5b, respectively. The loci imply multiple oscillations for the polarization evolution, as is confirmed later by the approximate analytical solutions in Sect. 2.2.2.
@@
Figure 5: For Example (iii): the evolution of the Stokes vector $\mathbf{S}\left(s\right)=\left(S_{1}\left(s\right),S_{2}\left(s\right),S_{3}\left(s\right)\right)$, $0\leq s\leq 20r_{*}\left(\approx 2\times 10^{7}\,\mathrm{cm}\right)$, for the X-ray emissions from the pulsar with $B_{\mathrm{max}}\approx 5.0\times 10^{13}\,\mathrm{G}$ and $\Omega_{\mathrm{p}}\approx 19.6\,\mathrm{Hz}$ ($P\approx 0.32\,\mathrm{s}$).
(a) Example (i)
(b) Example (ii)
(c) Example (iii)
Figure 6: Representations of the Stokes vectors from Examples (i)-(iii) on the Poincaré sphere. The loci imply patterns of the polarization evolution in terms of oscillation.
With regard to the adiabatic evolution condition as mentioned in Refs. Heyl2000 ; Heyl2003highenergy ; Heyl2018Strongly , we carefully examine our results presented in Figs. 3-5 to see what interpretations the condition leads to. Solving the condition $\left|k\mathbf{\hat{\Omega}}\left(\mathrm{d}\ln|k\mathbf{\hat{\Omega}}|/\mathrm{d}s\right)^{-1}\right|\gtrsim 0.05$ for $s$ yields $s_{\mathrm{PL}1}\lesssim s\lesssim s_{\mathrm{PL}2}$, [^10] Note that our $k\mathbf{\hat{\Omega}}$ is equivalent to the birefringent vector as defined in Refs. Heyl2000 ; Heyl2003highenergy ; Heyl2018Strongly and that we set the condition value to $0.05$ rather than $0.5$ as in the references. where $s_{\mathrm{PL}1[2]}$ refers to the lower [upper] bound for the ‘polarization limiting’ distance as measured from the emission point. Using this, one can check out the following:
(1) for $6.2\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 1.8\times 10^{6}\,\mathrm{cm}$ in Fig. 3,
(2) for $3.1\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 6.2\times 10^{6}\,\mathrm{cm}$ in Fig. 4,
(3) for $1.1\times 10^{5}\,\mathrm{cm}\lesssim s\lesssim 1.9\times 10^{7}\,\mathrm{cm}$ in Fig. 5,
our Stokes vector evolves evidently; otherwise, it freezes.